OpenStudy (anonymous):

What does momentum actually mean? (i mean ,does it not even have a definition) i refererd about 25 books but all i could get is --product of mass and velocity--if it cant be defined then how is it said that rate of change of this very momentum is force.............

OpenStudy (anonymous):

momentum=mass*change in velocity and acceleration=change in velocity/time since Force= mass * acc. so Force=mass*change in velocity/time Therefore; Force= momentum/time hence; Force = rate of change of momentum

OpenStudy (anonymous):

excuse meeee!!!!!! i asked the definition ........

OpenStudy (anonymous):

lol i dont know either exact definition of momentum

OpenStudy (anonymous):

"-if it cant be defined then how is it said that rate of change of this very momentum is force' I explained this part

OpenStudy (anonymous):

even my 7th class book has that stupid derivation

OpenStudy (anonymous):

This 'stupid' derivation answers your question if you try to understand it.

OpenStudy (anonymous):

I can tell you something which will help you understand it. 1) A body at rest will require a force to start moving, the more force you apply the more it will accelerate.. now you will have to apply large force to stop it.. this tendency of a body to get in rest after being in motion by again applying force in opposite direction is its momentum.. greater the velocity the greater force you have to apply to stop it.. I hope its clear I didn't include mass bcoz it's constant for a given body.

OpenStudy (anonymous):

momentum is simply the motion contained in a body.

OpenStudy (anonymous):

lols correct

OpenStudy (anonymous):

Momentum can be defined as "mass in motion." All objects have mass; so if an object is moving, then it has momentum - it has its mass in motion.

OpenStudy (anonymous):

It means the mass is on the move and is going to take some effort to stop. A team that has a lot of momentum is really on the move and is going to be hard to stop. Think about an impact. If the mass is small but the speed is really fast, then the momentum (imagine the impact) to you is really big (i.e. bullet). However if the speed is small but the mass is really big, you can get really hurt as well. In this picture is momentum.

OpenStudy (anonymous):

The amount of momentum that an object has is dependent upon two variables: how much stuff is moving and how fast the stuff is moving. That is why momentum = m*v

OpenStudy (anonymous):

Momentum of a mass is the instantaneous variation of its kinetic energy when we change its velocity

OpenStudy (anonymous):

This definition leads you to the Kinetic Energy:\[mv=dE_k/dv\] and \[dE_k=mvdv \rightarrow E_k=\int\limits mvdv=1/2mv^2\]

OpenStudy (anonymous):

@CarlosGP What you are telling him is useless since he do not understand any of it. I think your answer is intended to an examiner/teacher. If you are going to teach someone you have to explain in terms that they already have. What you're doing is just confusing him.

OpenStudy (anonymous):

@student92 1.-I have not heard him saying he does not understand any of it 2.-The answer is a physical concept, there are no physical concepts valid for teachers/examiners and other concepts valid for pupils 3.-If he had asked you what is velocity you would have said it is the variation of the location vs time but if I say momentum is the variation of Kinetic Energy vs speed then it is confusing 4.-This is blog to give the answers we know and the answers we can or we consider are best for a concept. There is freedom to ask, and there must be freedom to answer whenever code of conduct is respected. I appreciate comments if the answer is wrong or code of conduct is broken but comments because a third party thinks my answer is confusing can be perfectly made with a private message. 5.-I like your explanation about momentum

OpenStudy (anonymous):

both of u were good enough ,and can u tell y the proportionality const in it is 1............